OpenStudy (anonymous):

how to solve 2(2x+3)-10<6(x-2) ?

OpenStudy (anonymous):

move all terms with x to the left

OpenStudy (anonymous):

2(2x+3)-10<6(x-2) 4x+6-10<6x-12

OpenStudy (anonymous):

that the initial step

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

i dont kno wat to do after (-2x)<(-8)

OpenStudy (anonymous):

oh. multiply both sides by (-1)

OpenStudy (anonymous):

and this will change the sign

OpenStudy (anonymous):

< becomes >

OpenStudy (anonymous):

In this case: (-2x)<(-8) (-2x)*(-1)>(-8)*(-1) 2x>8 x>4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

but my buk ans says x<4

OpenStudy (anonymous):

2(2x+3)-10<6(x-2) 4x+6-10<6X-12 4x-4<6x-12 -4+12<2x 8<2x 4<x hope it helped a bit

OpenStudy (anonymous):

And 4<x is x>4

OpenStudy (anonymous):

So I guess your book is wrong this time @shrinifores

OpenStudy (anonymous):

As books cannot be right every time

OpenStudy (anonymous):

i'm confused wid this step 1) -2x<-8 2x>8 x>4 2) -2x<-8 x<-8/-2 x<4

OpenStudy (anonymous):

1) -2x<-8 2x>8 x>4

OpenStudy (anonymous):

This is correct

OpenStudy (anonymous):

do u want me to explain. why?

OpenStudy (anonymous):

yes, pls do...

OpenStudy (anonymous):

At first, -1>-2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so if u multiply both sides by -1 then, -1*(-1)>-2*(-2) 1>4 which is false

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

so there the sign changes... when we multily both sides by (-1) and divide both sides by (-1)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

but wat if we do this way.. -2x<-8 x<-8/-2 x<4

OpenStudy (anonymous):

Actually, there u jumped some few steps

OpenStudy (anonymous):

if u do it step by step... then u will see

OpenStudy (anonymous):

can u tel ?

OpenStudy (anonymous):

-2x<-8 -2x/2 <-8/2 -x<-4 -x*(-1)>-4*(-1) x>4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yeah now i get it... i cant solve as i do in an equation...

OpenStudy (anonymous):

YEAh....

OpenStudy (anonymous):

thank you so much....

OpenStudy (anonymous):

2(2x+3)-10<6(x-2) 4x+6 -10<6x-12 Disribute the 2(2x+3) by mulitplying 2*2=4. Add it on to the x, so 4x.Then 2*3=6 . Also, distribute the other side the same way 6*1x=6x and 6*-2=-12 4x-4<6x-12 -4x -4x Now subtract the 4x from both sides. -4<2x-12 +12 +12 Add the 12 to both sides 8<2x /2 /2 divide both side by 2 to get x by itself. 4<x This is the final answer.

OpenStudy (anonymous):

WELCOME

OpenStudy (anonymous):

^Just a simplified answer, because that's how my teacher taught it to us.

OpenStudy (anonymous):

i think i'm lacking in the basics...

OpenStudy (anonymous):

It's ok, and don't forget to flip the sign if you divide or multiply by a negative number.

OpenStudy (anonymous):

okie...

OpenStudy (anonymous):

BEST OF LUCK @shrinifores

OpenStudy (anonymous):

thnQ